Project

General

Profile

Cycle 2017/02 Startup discussion I

Highlight any major issues which would either prevent us starting up on time, or would
have an impact on performace.
Indicate if these issues can be fixed by Sept 19th and, if so, what resources are needed

  • Cycle 2017/01 Issues
  • Beamline
    • Conventional Magnets: no issues
    • Water system: need a week of running before needed for conductivity to recover. Turn on first week in Sept.
    • Beamstop: need updated documentation posting and BLOCs refresher.
    • Decay Solenoid: concerns - need input from Mark Tucker
    • Target: S19 is OK, test frame this week. R78 validation of S3 is problematic - slow progress. May then transfer big laser and newer control PC to production target.
    • Diffuser: tested during SSU ramp, OK. Proton Absorbers need inspection this week.
  • Detectors
    • TOFs/Ckov/KL: no issues
    • EMR: had a failing module in Feb./Mar. run - swapped in spare so none remaining.
    • Tracker: Did pump-purges during shutdown and fixed cryo 1 leak. Cooldown went much faster. Have run calibration: overall 10% dead channels up from 6%, cryo 1 at 26%. Have a proposed waveguide swapping 26->13% on cryo 1, 10->6% overall. Are analysing how to optimise dead individual channels vs. "holes" and separately looking for electronics-side issues; analyses will take a couple of days each. Decide on plan next week and report next Ops Meeting. Waveguide changes should take a day but verifying the channel mapping then takes a week.
  • CC magnets
    • SSU/SSD: ramped to nominal, missing QPS signal on SSU M2. Issue with connector needs to be fixed before Magnet Ramp-Up.
    • Focus coil
  • LH2: sign-off tour on 31st, then test fill-level-empty cycle week of 4th?
  • Hall
    • PPS
    • Air conditioning: unit 3 off, 4 has filter warning, 1 off and known to be weak.
  • Alarms and Monitoring: see issues. New thing is the masking-off procedure for persistent alarms; in the meantime new alarms must be investigated and solved. Final pass through persistent alarm list and disable unwanted ones: 18th Sept. SBoyd will re-jig shift handover checklist to ensure things are recorded.
  • DAQ and Online: ISIS RF signal is needed in RR1. SBoyd to check with ISIS and get Craig to confirm it exists when we think it's live
    • Run Control: works at the moment as long as people don't rush. Need to get LH2 Absorber information - level, etc. - from EPICS into CDB. Durga would prefer to do this by post-hoc correction rather than trying to modify Run Control.
    • CDB
    • OnRec: will need to upgrade to new MAUS version.
    • Data Transfer: will need to watch storage space on miceraid5 and micestore. DR to notify when Step 1 can be deleted from latter
    • CDB Viewer crashes browser windows in MLCR. Need to downgrade Firefoxes? PFranchini will check. Re-check all machines during start-up.
  • Staffing
    • Shifter training: arranged similarly, need to define LH2 requirements (SBoyd) and repeat for everyone
    • BLOCs: rota has gaps near start and in middle - 3 stints. Should be on day shift if doubling up.

Next meeting 4th September.

S19.pdf - S19 Target plots. Latest data is included however data from when the computer broke is omitted (Sep-Dec 2016) for now hence why it shows less than 5 million. (68 KB) Langlands, Joe, 23 August 2017 10:29

Irony.png - The Internet is getting too smart... (245 KB) Nebrensky, Henry, 01 September 2017 15:34